# Privacy Preserving Mechanisms for Coordinating Airspace Usage in Advanced Air Mobility

### v1

@@ Abstract
Advanced Air Mobility (AAM) operations are expected to transform air transportation while challenging current air traffic management practices. By introducing a novel market-based mechanism, we address the problem of the on-demand allocation of capacity-constrained airspace to AAM vehicles with heterogeneous and private valuations. We model airspace and air infrastructure as a collection of contiguous regions (or sectors) with constraints on the number of vehicles that simultaneously enter, stay, or exit each region. Vehicles request access to the airspace with trajectories spanning multiple regions at different points in time. We use the graph structure of our airspace model to formulate the allocation problem as a path allocation problem on a time-extended graph. To ensure that the cost information of AAM vehicles remains private, we introduce a novel mechanism that allocates each vehicle a budget of “air-credits” (an artificial currency) and anonymously charges prices for traversing the edges of the time-extended graph. We seek to compute a competitive equilibrium that ensures that: (i) capacity constraints are satisfied, (ii) a strictly positive resource price implies that the sector capacity is fully utilized, and (iii) the allocation is integral and optimal for each AAM vehicle given current prices, without requiring access to individual vehicle utilities. However, a competitive equilibrium with integral allocations may not always exist. We provide sufficient conditions for the existence and computation of a fractional-competitive equilibrium, where allocations can be fractional. Building on these theoretical insights, we propose a distributed, iterative, two-step algorithm that: 1) computes a fractional competitive equilibrium, and 2) derives an integral allocation from this equilibrium. We validate the effectiveness of our approach in allocating trajectories for two emerging urban air mobility services: drone delivery and air taxis.
@@ 1 Introduction
The emergence of advanced air mobility (AAM) operations, including urban air mobility (UAM) and unmanned aerial vehicles (UAVs), is expected to transform the landscape of the air transportation system. These new aerial platforms can provide air taxi services that better connect rural and suburban communities with urban centers, facilitate package and medical delivery services, and support infrastructure and public safety (Guo et al., 2024). However, these new technologies bring new challenges concerning air traffic management. Current estimates of the density, type, and number of these new flights have led the Federal Aviation Administration (FAA) in the United States to declare that “existing Air Traffic Management (ATM) system infrastructure and associated resources cannot cost-effectively scale to deliver services” (Administration, 2022). While current systems focus primarily on fixed-wing aircraft, scheduled flight operations, and airport infrastructure, AAM includes novel electric vertical take-off and landing (eVTOL) aircraft and UAVs that will fly on demand, which creates challenges for scheduling flights. Additionally, like current aerial operations (Ball et al., 2018, 2020), AAM operations will have heterogeneous valuations for timely airspace access (while a passenger air taxi may be traveling on a strict schedule, a regional cargo flight could be delayed with less issue (Skorup, 2019; Seuken et al., 2022)), yet it is extremely important for these valuations to remain private because they can contain important business or personal information. These characteristics have led the FAA and other aviation agencies to call for novel air traffic management tools and strategies, developed by third-party Service Providers (SPs), to support AAM aircraft and use cases. These tools must work in conjunction with existing Air Navigation Service Providers (ANSPs, e.g., the FAA) to safely and efficiently actualize new air transport opportunities (Administration, 2022; Federal Aviation Administration, 2023; European Organisation for the Safety of Air Navigation (2022), EUROCONTROL).
We propose an algorithm that schedules AAM vehicles with heterogeneous, private valuations requesting access to airspace in an on-demand manner. While ideas from conventional air traffic management (Bertsimas and Patterson, 1998, 2000; Bertsimas et al., 2011; Roy and Tomlin, 2007; Odoni, 1987) can accommodate private and heterogeneous valuations from flight operators, they are not designed for the dynamic and adaptable nature of AAM operations (Skorup, 2019; Seuken et al., 2022). This work aims to answer the following question:
@@
Can we design an efficient mechanism that allocates desired airspace resources to AAM vehicles with heterogeneous private valuations without requiring them to disclose their valuations to anyone?
@@
To address this, we model the airspace as a set of contiguous regions, each with specific capacity constraints on the number of AAM vehicles (modeled as eVTOLs) that can arrive, depart, or stay at any given time (see Fig. 1). Vehicles submit requests for airspace access through a menu of time-trajectories (or air corridors), detailing which regions they will occupy at different time steps. We frame the allocation of these time-trajectories within capacity-constrained airspace as a path allocation problem on a time-extended graph (Definition 2.1), where all capacity constraints are represented as constraints on the graph’s edges.
@@
We introduce a new mechanism with artificial currency that is implemented in a receding horizon manner. In every round of this approach, the SP allocates “air-credits” (the artificial currency) to each AAM vehicle requesting access to airspace in that round and charges an anonymous price (in air credits) for the use of various airspace regions. Based on these prices, each vehicle calculates its most preferred bundle of regions on the time-extended graph that maximizes its valuation while adhering to its budget constraint and some additional individual constraints to ensure that the set of bundles forms a feasible path.
@@
The goal of the SP is to design prices and allocate airspace regions efficiently and safely, guided by the following desiderata: $(i)$ Given the price vector, the allocation by the SP should be optimal for every AAM vehicle, maximizing each vehicle’s private valuation subject to budget constraints; $(ii)$ The capacity constraints of the airspace must be respected; $(iii)$ Prices must be non-negative, and if they are strictly positive, the capacity of each airspace region must equal its demand. A tuple of allocation and prices that satisfy these requirements is referred to as a competitive equilibrium in economics, which may not always exist (Budish, 2011). However, inspired by the literature on Fisher markets under linear constraints (Jalota et al., 2023), we demonstrate that a fractional competitive equilibrium — a relaxation of the competitive equilibrium that allows fractional allocations — always exists (Proposition 3.3). Furthermore, we show that the prices at a fractional-competitive equilibrium can be computed as an optimal dual multiplier of a ”budget-adjusted welfare problem” (see (3)), which is a convex optimization problem (Lemma 3.4). Notably, the budget adjustment for each vehicle is determined by the optimal dual multiplier associated with its individual constraints. As a result, computing a fractional-competitive equilibrium reduces to calculating a fixed point (Proposition 3.5).
@@
Building on these theoretical insights, we propose a two-step algorithmic procedure to allocate AAM vehicles to airspace. In the first step, we develop a two-loop algorithm to compute the fractional-competitive equilibrium without requiring information about the private valuations of vehicles. Specifically, this step involves solving the fixed point problem stated in Proposition 3.5 via a two-loop algorithm (see Algorithm 1) that mimics fixed point iteration. The inner loop solves a reformulated budget-adjusted welfare problem in a distributed manner using an Alternating Direction Method of Multipliers (ADMM) update, ensuring that AAM vehicles do not have to share their valuations with anyone. The outer loop then updates the budget adjustment parameter using the most recent value of the dual multiplier corresponding to the individual constraints of the AAM vehicles.
@@
In the second step, we derive an integral allocation from the fractional-competitive equilibrium obtained in the first step, keeping the prices unchanged. We rank the vehicles based on the fractional allocation they received for their most desired resource in the first step. The social planner then allocates resources to the vehicles sequentially according to this ranking, updating the remaining capacity after each allocation (Algorithm 2).
@@
Algorithm 1 can be viewed as feedback between the learning algorithm used by AAM vehicles and the SP. The AAM vehicles use (anonymized) signals from the SP, such as prices and other parameters, to adjust their demand for resources. The SP, in turn, uses the updated demand for resources from each AAM vehicle to adjust the prices and related variables. Notably, this two-step process (Algorithms 1-2) does not require vehicles to disclose their private valuations to either the SP or to other vehicles (Remark 4.2). To validate the effectiveness of our approach, we explore two upcoming applications. First, we analyze drone-based package delivery using a dataset of drone trajectories generated with realistic physical models by Airbus over the city of Toulouse, France. Second, we investigate the scheduling problem for electric air taxis on a hypothesized air traffic network in Northern California, United States.
@@ 1.1 Related Works
Market mechanisms for airspace management.
Su et al. (Su et al., 2024) utilized a generalized Vickrey–Clarke–Groves (VCG) auction for AAM resource management taking into account social welfare, safety, and congestion while considering proportional fairness by weighting different fleet operators accordingly. We adopt a similar time-extended graph model from their work to represent airspace with capacity constraints. However, when agents have budget constraints, as in our case, truthful bidding is no longer a dominant strategy (Ausubel and Milgrom, 2004), which limits the applicability of VCG-based mechanisms. Furthermore, the requirement for truthful bidding poses privacy concerns for vehicles as they have to reveal their exact valuations through bids. Other proposed approaches include second-price auctions combined with congestion management algorithms (Qin and Balakrishnan, 2022; Chin et al., 2023) and combinatorial auctions (Leet and Morris, ). Unlike our approach, these related works only considered unit-capacity regions.
Balakrishnan and Chandran’s work (Balakrishnan and Chandran, 2017) introduced a column generation algorithm that dynamically updates prices to respect capacity constraints, approximating a competitive equilibrium. Their approach, however, requires aircraft to report both their desired trajectory and the associated benefits. In contrast, our ADMM-based approach simplifies the process by requiring aircraft to report only their desired trajectories in response to announced prices. While the former approach may benefit from computational efficiency due to the additional information available to the central planner, our approach prioritizes privacy by limiting the amount of information shared. Another key distinguishing feature of our work, compared to existing market mechanisms for AAM, is the use of artificial currency. While, as noted by (Seuken et al., 2022), money can effectively elicit preferences, mechanisms involving monetary transactions may disproportionately favor operators with greater financial resources. Our approach mitigates this issue by introducing a system of artificial currency that helps ensure fairness.
Connections with other economic mechanisms.
Our airspace allocation formulation incorporates four essential features: complementarities, indivisibilities, the absence of monetary transfers, and quasilinear preferences. In settings with substitutable and divisible goods, the classical concept of Competitive Equilibrium from Equal Incomes (CEEI) ensures efficient and fair outcomes by endowing agents with an artificial currency that holds no value outside the market (Varian, 1974). However, in cases involving complementary and indivisible goods, a competitive equilibrium may not exist. A common approach in such cases is to seek an approximate CEEI (A-CEEI), which preserves efficiency, fairness, and incentive properties while being supported by computationally efficient algorithms tested in real-world applications like course allocation (Budish, 2011; Budish et al., 2023).
Although we adopt the idea of artificial currency, we deviate from this model by allowing agents to save currency for future use. When the saved budget influences the agent’s utility in a quasilinear form, our model more closely aligns with combinatorial auction environments, where monetary transfers play a role in eliciting agents’ preferences.
@@
Combinatorial auctions allow participants to bid on combinations of items rather than individual items (Cramton et al., 2006a). Given the inherent exponential complexity of these auctions, there is no universal format that applies in all settings. Below, we highlight formats particularly suited to scenarios with budget constraints and strong complementarities. For a more comprehensive list, refer to (Cramton et al., 2006b, Table 2.2). In environments with budget constraints, dynamic auction formats often outperform static mechanisms like VCG auctions. These formats allow agents to observe current prices and adjust their bids according to their budget constraints before final submission. Examples include the Simultaneous Ascending Auction, the Ascending Proxy Auction, and the Clock-Proxy Auction (Ausubel and Cramton, 2004; Cramton et al., 2006b). These dynamic formats have an additional advantage: bids are submitted incrementally, reducing the risk of privacy violations and preventing the exponential growth of bundle combinations. However, they are still susceptible to computational complexity issues, which can even be exacerbated due to the inherent nature of combinatorial auctions. In our approach, we model complementarities using linear equality constraints and leverage recent advances in Fisher markets with linear constraints to develop an algorithmic approach that only needs to solve convex optimization problems.
Particularly, we compute a fractional-competitive equilibrium, which provides a relaxed solution that is easier to compute. From this fractional-competitive equilibrium, we then derive an integral allocation that satisfies the capacity constraints in the problem.
@@ 1.2 Organization
The article is organized as follows: In Section 2, we introduce the model of airspace management studied in this paper. In Section 3, we describe the proposed market mechanism with artificial currency and the theoretical results on fractional-competitive equilibrium. Section 4 presents the algorithmic procedure to compute an approximate market mechanism. In Section 5, we validate the performance of our mechanism using a drone delivery dataset generated based on a real drone dynamics model from Airbus. In Section 6, we study an additional AAM scenario of vertiport reservation for air-taxi services in Northern California. We conclude this study in Section 7 and state some interesting directions for future work. Proofs and additional supplementary material are provided in the appendix. In Section A, we present a simple example to illustrate the time-extended graph and constraints (2b)-(2c). In Section B, we provide proofs for all theoretical results discussed in this paper. Finally, in Section C, we consolidate all important notations used in this work in the form of a table.

### v2

@@ Abstract
Advanced Air Mobility (AAM) operations are expected to transform air transportation while challenging current air traffic management practices. By introducing a novel market-based mechanism, we address the problem of on-demand allocation of capacity-constrained airspace to AAM vehicles with heterogeneous and private valuations. We model airspace and air infrastructure as a collection of contiguous regions (or sectors) with constraints on the number of vehicles that simultaneously enter, stay, or exit each region. Vehicles request access to airspace with trajectories spanning multiple regions at different times. We use the graph structure of our airspace model to formulate the allocation problem as a path allocation problem on a time-extended graph. To ensure that the cost information of AAM vehicles remains private, we introduce a novel mechanism that allocates each vehicle a budget of “air-credits” (an artificial currency) and anonymously charges prices for traversing the edges of the time-extended graph. We seek to compute a competitive equilibrium that ensures that: (i) capacity constraints are satisfied, (ii) a strictly positive resource price implies that the sector capacity is fully utilized, and (iii) the allocation is integral and optimal for each AAM vehicle given current prices, without requiring access to individual vehicle utilities. However, a competitive equilibrium with integral allocations may not always exist. We provide sufficient conditions for the existence and computation of a fractional-competitive equilibrium, where allocations can be fractional. Building on these theoretical insights, we propose a distributed, iterative, two-step algorithm that: 1) computes a fractional competitive equilibrium, and 2) derives an integral allocation from this equilibrium. We validate the effectiveness of our approach in allocating trajectories for the emerging urban air mobility service of drone delivery.
@@ 1 Introduction
The emergence of advanced air mobility (AAM) operations, including urban air mobility (UAM) and unmanned aerial vehicles (UAVs), is expected to transform the landscape of the air transportation system. These new aerial platforms can provide air taxi services that better connect rural and suburban communities with urban centers, facilitate package and medical deliveries, and support infrastructure and public safety (Guo et al., 2024).
Integrating advanced air mobility (AAM) into existing air traffic management (ATM) systems presents complex and unresolved challenges. Projections of AAM traffic density and operational complexity have raised concerns about the scalability of traditional ATM infrastructure. The Federal Aviation Administration (FAA) has acknowledged these challenges (Administration, 2022; Federal Aviation Administration, 2023), stating:
“Given the number, type, and duration of Unmanned Aircraft System (UAS) operations envisioned, the existing Air Traffic Management (ATM) System infrastructure and associated resources cannot cost-effectively scale to deliver services for UAS.” — FAA UAS/UTM Con-Ops ((Administration, 2022))
The limitations of conventional ATM approaches become evident when examining their design principles. Existing systems (Bertsimas and Patterson, 1998, 2000; Bertsimas et al., 2011; Roy and Tomlin, 2007; Odoni, 1987; Ball et al., 2018, 2020) are built to manage fixed-wing aircraft operating between established airports, with flight schedules planned weeks or months in advance to minimize overall delays. In contrast, AAM introduces a fundamentally different paradigm, with a high volume of electric Vertical Take-Off and Landing (eVTOL) aircraft and UAVs operating on demand and pursuing diverse objectives. These vehicles will not only travel between fixed vertiports but also serve ad-hoc destinations, such as residential areas for package deliveries, further straining legacy ATM systems. To address this challenge, the FAA (Administration, 2022; Federal Aviation Administration, 2023) and other Air Navigation Service Providers (ANSPs) worldwide (European Organisation for the Safety of Air Navigation (2022), EUROCONTROL) have stated that daily traffic management for AAM operations will be delegated to third-party service providers (SPs). These providers will coordinate directly with AAM vehicles to allocate airspace efficiently and safely, reducing reliance on the FAA.
Beyond operational challenges, the introduction of third-party SPs in AAM systems raises additional concerns, particularly regarding privacy. A key issue arises from the heterogeneous private valuations of AAM vehicles, which can vary significantly depending on their specific use cases. For instance, a passenger air taxi may have strict scheduling constraints, whereas a regional cargo flight might be more flexible with delays (Skorup, 2019; Seuken et al., 2022). This variability in preferences, coupled with the sensitivity of business or personal data, makes AAM operators reluctant to disclose their private valuation information to SPs.
Against this backdrop, this work aims to answer the following question:
@@
How can SPs allocate capacity-constrained airspace resources to dynamically arriving AAM vehicles with heterogeneous private valuations in a way that allows AAM vehicles to achieve an (approximately) optimal allocation based on their valuations, without requiring them to disclose this private information to the SPs?
We address this question by focusing on a single service provider managing access to a specific region of airspace, as even under a single SP, efficiently allocating airspace to AAM vehicles without requiring their private information remains an open problem.
@@
We model the airspace as a set of contiguous regions, each with specific capacity constraints on the number of AAM vehicles (modeled as eVTOLs) that can arrive, depart, or remain in a given region at any time (see Fig. 1). Vehicles submit requests for airspace access through a menu of feasible (discrete-time) time-trajectories (or air corridors), where each time-trajectory corresponds to a sequence of tuples specifying the time step and the sector they wish to occupy at that time. We formulate the allocation of these time-trajectories within capacity-constrained airspace as a path allocation problem on a time-extended graph (Definition 2.1), where all capacity constraints are represented as constraints on the graph’s edges.
@@
To prevent the monopolization of airspace by major players, we introduce an artificial currency-based auction mechanism (Skorup, 2019). Furthermore, to accommodate the dynamically arriving requests of AAM vehicles, we propose implementing this auction mechanism in a receding-horizon manner (see Section 3). This approach periodically collects AAM vehicle requests and determines allocations using the proposed auction mechanism.
@@
In each of these auction mechanisms, the SP allocates “air-credits” (the artificial currency) to each AAM vehicle requesting airspace access and charges an anonymous price (in air-credits) for using different airspace regions (i.e., resources on the time-extended graph). Based on these prices, each vehicle selects its most preferred time-trajectory on the time-extended graph, maximizing its valuation while adhering to its budget constraint. The SP’s objective is to design prices and allocate airspace efficiently and safely, guided by the following desiderata: $(i)$ Given the price vector, the SP’s allocation should be optimal for every AAM vehicle, ensuring that each vehicle maximizes its private valuation subject to budget constraints; $(ii)$ The capacity constraints of the airspace must be respected; $(iii)$ Prices must be nonnegative, and if strictly positive, the capacity of each airspace region must match its demand. An allocation-price tuple satisfying these conditions is known as a competitive equilibrium in economics, which may not always exist (Budish, 2011). However, drawing inspiration from Fisher markets under linear constraints (Jalota et al., 2023), we establish the existence of a fractional competitive equilibrium—a relaxation of the competitive equilibrium that permits fractional allocations (Proposition 4.3). Moreover, we demonstrate that the prices at a fractional competitive equilibrium can be computed as the optimal dual multipliers of a budget-adjusted welfare problem (see (3)), which is a convex optimization problem (Lemma 4.4). Notably, the budget adjustment for each vehicle is determined by the optimal dual multiplier associated with a linear constraint of this optimization problem. Consequently, computing a fractional competitive equilibrium reduces to solving a fixed-point problem (Proposition 4.5).
@@
Building on these theoretical insights, we propose a two-step algorithmic procedure for allocating AAM vehicles to airspace. In the first step, we develop a two-loop algorithm to compute the fractional competitive equilibrium without requiring information about the vehicles’ private valuations. Specifically, this step involves solving the fixed-point problem stated in Proposition 4.5 using a two-loop algorithm (see Algorithm 1) that mimics fixed-point iteration. The inner loop solves a reformulated budget-adjusted welfare problem (cf. (4)) in a distributed manner using the Alternating Direction Method of Multipliers (ADMM) (see Appendix C for a review of ADMM). This ensures that AAM vehicles do not need to share their valuations with SP and other AAM vehicles. The outer loop then updates the budget adjustment parameter using the latest value of the dual multiplier associated with each AAM vehicle’s individual constraints.
Algorithm 1 can be interpreted as an online learning process, where AAM vehicles iteratively refine their trajectory choices based on anonymized market signals—such as expected demand and prices—while the SP dynamically adjusts these signals based on observed demand by AAM vehicles. This enables the SP to estimate equilibrium prices without access to the private valuations of AAM vehicles, while vehicles deconflict through indirect price-mediated coordination, obviating the need for direct trajectory or valuation sharing. This mechanism aligns with online learning of market mechanisms (cf. (Kleinberg and Leighton, 2003; Li et al., 2020; Bistritz and Bambos, 2021)), as both the agents and the SP update their strategies sequentially based on observed feedback.
@@
In the second step (i.e., Algorithm 2), we derive an integral allocation from the fractional competitive equilibrium obtained in the first step while keeping the prices unchanged. We rank the vehicles according to the fractional allocation they received for their most desired resource in the first step. The SP then allocates resources to the vehicles sequentially according to this ranking, updating the remaining capacity after each allocation (Algorithm 2). Importantly, in both Algorithms 1-2, the SP requires only information on resource demands, feasible time trajectories, and the most desired path of each AAM vehicle, without accessing any private valuation data. Likewise, individual AAM vehicles do not obtain information about the time trajectories or private valuations of other vehicles, ensuring that privacy is preserved throughout the process.
@@
To validate the effectiveness of our approach, we analyze drone-based package delivery using a dataset of drone trajectories generated with realistic physical models by Airbus over the city of Toulouse, France. A further study on the scheduling problem for electric air taxis on a hypothesized air traffic network in Northern California, United States can be found in Appendix D.
@@ 1.1 Related Works
The literature on market mechanisms for airspace management in Advanced Air Mobility (AAM) remains relatively limited (Seuken et al., 2022). Some recent works (Qin and Balakrishnan, 2022; Chin et al., 2023b) have explored airspace management through second-price auctions combined with congestion management algorithms, as well as combinatorial auctions (Leet and Morris, ). However, these approaches are restricted to unit-capacity regions, limiting their applicability to real-world AAM systems that require flexible allocation across multiple capacity-constrained airspace sectors. Su et al. (Su et al., 2024) addressed these limitations by utilizing a generalized Vickrey–Clarke–Groves (VCG) auction for AAM resource management, incorporating considerations of social welfare, safety, and congestion. Their approach ensures proportional fairness by optimizing a social cost function based on the weighted utilities of all fleet operators. A key design goal in their paper is to elicit truthful preferences from AAM vehicles to achieve efficient airspace allocation. However, this reliance on truthful bidding raises privacy concerns, as vehicles must disclose their exact valuations through bids, potentially exposing sensitive operational information to competitors or regulators. Balakrishnan and Chandran (Balakrishnan and Chandran, 2017) proposed a column generation algorithm that iteratively updates prices to determine an allocation that satisfies capacity constraints. However, their method relies on vehicles reporting their private valuations, which may raise privacy concerns. In contrast, our approach eliminates the need for AAM vehicles to disclose their private valuations, enhancing privacy while still achieving efficient allocation.
A distinguishing feature of our approach compared to existing market mechanisms for AAM is the use of artificial currency. While monetary transactions are effective in eliciting preferences, as noted by (Seuken et al., 2022), they can disproportionately advantage operators with greater financial resources. Our approach mitigates this issue by introducing a system of artificial currency that helps ensure fairness. The idea of using artificial currency for fair and efficient resource allocation has been extensively studied in economics, beginning with (Varian, 1974). These works typically consider environments where agents are endowed with artificial currency that holds no value outside the market. However, most of these studies focus on the allocation of substitutable goods, whereas our setting requires agents to select multiple goods over a time-extended network, leading to complementarities rather than substitutability. Artificial currency mechanisms that handle complementarities have been studied in (Budish, 2011; Budish et al., 2023). The key distinction of our work is that we allow agents to save currency for future use, and the saved budget influences the agent’s utility in a quasi-linear manner. This feature aligns our model more closely with combinatorial auction environments.
@@
Combinatorial auctions enable participants to bid on bundles of items rather than individual items (Cramton et al., 2006a). Given the exponential complexity inherent in these auctions, there is no single format that universally applies across all settings. In environments with budget constraints, iterative auction formats—such as the Simultaneous Ascending Auction, the Ascending Proxy Auction, and the Clock-Proxy Auction (Ausubel and Cramton, 2004; Cramton et al., 2006b)—are particularly relevant. These formats allow agents to observe current prices and iteratively adjust their bids within their budget constraints before final submission. Additionally, they offer a privacy advantage by enabling incremental bid submission, thereby reducing exposure of full preference information. However, these mechanisms may suffer in efficiency in combinatorial auction settings with strong complementarities. In our approach, we propose a new way to model complementarities using linear equality constraints and leverage recent advances in Fisher markets with linear constraints to determine allocation. In particular, we compute a fractional-competitive equilibrium, which provides a relaxed solution that is easier to compute. From this fractional-competitive equilibrium, we then derive an integral allocation that satisfies the capacity constraints in the problem.
@@ 1.2 Organization
The article is organized as follows: In Section 2 we introduce the model of airspace management studied in this paper. In Section 3 we present a high-level overview of our approach, which implements an artificial-currency-based auction mechanism in a receding horizon manner. Next, in Section 4 we formally present such an auction mechanism along with theoretical results on fractional-competitive equilibrium. In Section 5, we present the algorithmic procedure to compute an approximate market mechanism. In Section 6, we validate the performance of our mechanism using a drone delivery dataset generated based on a real drone dynamics model from Airbus. We discuss the limitations of our approach in Section 7, and conclude this study in Section 8 with some interesting directions for future work. Proofs and additional supplementary material are provided in the Appendix. In Appendix A, we present a simple example to illustrate the time-extended graph and constraints (2b)-(2c). In Appendix B, we provide proofs for all the theoretical results discussed in this paper. A detailed explanation of the ADMM formulation can be found in Appendix C. In Appendix D, we study an additional AAM scenario of vertiport reservation for air-taxi services in Northern California. Finally, in Appendix E, we consolidate important notations used in this work in a table.